If a customer receives a 15% discount on an $875.00 item, what will their final cost...

  1. If a customer receives a 15% discount on an $875.00 item, what will their final cost be? ________
  2. Your customer is looking for new flooring for their kitchen. The room is 14’ x 24’. 4 boxes of flooring will cover 100 sq. ft. How many boxes will be needed to cover the kitchen floor?
  1. A builder wants to purchase countertops that are 15’ long and 24” deep, plus an island countertop that is 48” x 36”, how much total square footage will be needed for this?
Answer #1

solution Discount = 15%. Initial cost = $875.00 final cost = Initial cost - Discount = (875-15 x 875 ) $ = $ 743.75 final cosDate Page - 12 inches total sauare for tre rootage needed = 15xQ ) + ( 4x 3) = (30+12) And - Ga Atd total syuare footage need